Solve the following system by substitution
y=2x+3
5x-2y+ -8

Found 2 solutions by 10301650, mananth:
Answer by 10301650(32)

y=2x+3
5x-2y -8
.
.
---: use the second given to find the answer for x
5x-2y-8
---: then substitute the variable y into 2x+3 (the 1st given)
5x-2(2x+3)-8=0
5x-4x+6-8=0
x-2=0
x=2 (ANSWER)
.
.
---: now lets solve for the y. use the 1st the 1st given
.
y=2x+3
y=2(2)+3
y=4+3
y=7 (ANSWER)

Answer by mananth(16946)

5 x -2 y = -8 .............1
1 y = 3 + 2 x ---------2
/ 1 =
y = 3 + 2 x
Substitute y in (1)
5 x -2 ( 3 + 2 x ) = -8
5 x -6 + -4 x = -8
5 x -4 x = -8 6
1 x = -2
/ 1
x= -2
Plug the value of x in (1)
5 x -2 y = -8
5 * -2 -2 y = -8
-10 -2 y = -8
-2 y = 2
/ -2
y= -1
m.ananth@hotmail.ca